OpenStudy (anonymous):

Solve and show work! (1/n^2) + (1/n) = (1/2n^2)

OpenStudy (dumbcow):

\[\frac{1+n}{n^{2}} = \frac{1}{2n^{2}}\] \[2n^{2} +2n^{3} = n^{2}\] \[2n^{3}+n^{2} = 0\] \[n^{2}(2n+1) = 0\] n = 0, -1/2

OpenStudy (dumbcow):

but 0 is not valid because you can't divide by 0